# HR-INR: Continuous Space-Time Video Super-Resolution via Event Camera

### v1

@@ 3.1 Temporal Pyramid Representation
Firstly, we represent the event stream $\boldsymbol{E}$ into a frame-like form that the network can process. To capture holistic motion, we partition all events during $[0,1]$ of the timeline into $M$ equal intervals using a voxel grid [62, 14], denotes as $\boldsymbol{E}_{v}$ with dimensions $M\times H\times W$. In practice, event representation methods like voxel grid [62, 14] and its extended structure, symmetric cumulative event representation [57], achieve time granularity by uniformly dividing time into intervals with resolutions of $1/M$.
@@
Our idea: The core of TPR is constructing a temporal pyramid where each successive layer’s duration is $1/r,(r>1)$ of the preceding one, leading to exponentially finer time granularity with additional layers. For instance, as illustrated in Fig. 2 (a), around any given time $t$, we define a surrounding time window $\Delta t$ and select an attenuation factor $r$. At the pyramid’s $L$-th level, the events are within the time span of $[t-\Delta t/r^{L},t+\Delta t/r^{L}]$. Each layer is further segmented into $M_{p}$ intervals, represented using a voxel grid. Accordingly, for an $L$-th layer, each layer contains $M_{p}$ moments within the TPR, and its finest time granularity, denoted by $\delta_{t}$, is as delineated in the Eq. 2:
@@
Therefore, for any time $t$, we construct the corresponding TPR $E_{p}$ with shape $L\times M_{p}\times H\times W$. We record the TPRs at all target timestamp as $\boldsymbol{E_{p}}=\{E_{p}\}_{i}^{N_{out}}$.
Discussion: The time granularity $\delta_{t}$ of TPR exponentially improves with the increase in layers, $L$. For an attenuation factor of $r=3$ and a goal to detect motions as brief as $1/1000$ of a second within a $1s$ window ($12\times\Delta t=$ 1), we require only $7$ TPR layers with each layer divided into $2$ intervals ($M_{p}=2$). Consequently, a TPR with dimensions $7\times 2\times H\times W$ suffices to discern motion down to $1/1000s$. Based on the above representation, we obtained $\boldsymbol{E}_{v}$, encapsulating holistic motion, and $\boldsymbol{E}_{p}$, which focuses on regional edges and motion.
@@ 3.2 Holistic-Regional Feature Extraction
This module aims to extract features from regional TPRs $\boldsymbol{E_{p}}$, and the frame $\boldsymbol{I_{in}}$ and the holistic events $\boldsymbol{E_{v}}$ for INR-based spatial-temporal decoding. Accordingly, we introduce: (1) the regional event feature extractor, $f_{re}$, for dynamic motion and edges detail capture. (2) the holistic event-frame feature extractor, $f_{he}$, for long-term motion dependencies modeling across time and space.
@@
Figure 3: Holistic event-frame feature extractor. The down-sample module will halve the resolution. The up-sample module will double the resolution. The encoder and decoder have the same structure as Swin-Transformer [37, 35, 16].
Regional Event Feature Extraction: The input of $f_{re}$ is TPR $E_{p}\in R^{L\times M_{p}\times H\times W}$ in each timestamp $t$. Given that $f_{re}$ is invoked $N_{out}$ times, its design needs to strike a balance between efficiency and the ability to model inter-level relationships to ensure that it can capture precise regional motion and edge details effectively. Initially, our method applies a convolution layer to increase the dimensions of the feature. Then, we use four Swin Transformer Encoder Blocks (STEB) [37, 35, 16] to model the relationship between different pyramid levels. The STEB enjoys a large view field and a multi-head attention mechanism that proficiently models distant dependencies and conveys edge information across various pyramid levels, effectively capturing short-term motion. Notably, STEB is optimized with fewer parameters, boosting computational efficiency.
Holistic Event-Frame Feature Extraction: The inputs of $f_{he}$ is $\boldsymbol{I_{in}}$ and $\boldsymbol{E_{v}}$, as shown in Fig. 3. First, a convolutional layer processes both frames and events to increase dimensions. Motivated by the events feature manifest between successive frames to provide inter-frame motion information. We adjusted the first dimension of the event feature to $N_{in}-1$ to account for inter-frame motion losses. We then incorporate the STEB to facilitate interactions across varied levels and spatial domains. To minimize computational overhead and expand the receptive field, we integrated a down-sampling module between STEBs, forming a multi-scale encoder. Each down-sampling iteration halves the resolution while maintaining the channel dimensions. After three iterations, the feature resolution reduces to $1/8$ of its initial size, enlarging the receptive field. We then employ a Swin Transformer Decoder Block [37, 35, 16] to fuse features at matching resolutions. Each of the first three STEBs is followed by an upsampling process, which doubles the resolution while maintaining the channel count. Ultimately, this process outputs the feature $F_{g}$. For more details, see Suppl. Mat..
Notably, to output $N_{out}$ frames, the holistic event-frame feature extractor $f_{he}$ is called once, capturing the comprehensive feature $F^{g}$. Subsequently, for each time $t$, $f_{re}$ extracts regional features $F_{t}^{l}$ from each TPR $E_{p}\in\boldsymbol{E_{p}}$. For each regional feature $F_{t}^{l}$, we use addition and Cov1×$\times 1$ 1 operation as fusion function $f_{fu}$ to fuse with holistic feature $F^{g}$ to obtain the output $R_{t}$. For each time $t$, the whole process can be described by Eq. 3, where $E_{p}\in\boldsymbol{E_{p}}$ is the TPR at the specific timestamp $t$.
$\begin{split}R_{t}=&f_{fu}\left(F^{g},F_{t}^{l}\right);F^{g}=f_{he}\left(\boldsymbol{I_{in}},\boldsymbol{E_{v}}\right);F_{t}^{l}=f_{re}\left(E_{p}\right)\end{split}$ (3)
@@
Figure 4: Temporal embedding. Given the input time $t\in[0,1]$, the output is the temporal attention $E_{t}$ derived from a two-layer MLP. (b) presents a visualization of the trained $E_{t}$ during $[0,1]$ on real-world dataset [61].
@@ 3.3 INR-based Spatial-Temporal Decoding
Spatial Embedding and Decoding: To upscale the temporal embedded features to any desired spatial scale, we utilize a similar approach to previous works [10, 8, 9]. We query the four nearest neighbors in the temporal embedded feature for each spatial coordinate and concatenate these with their distances for spatial embedding. A four-layer MLP decoder computes the RGB values, which are then aggregated through area-weighted interpolation for arbitrary-scale super-resolution. Similar to [38], we use the Charbonnier loss [30] as the fundamental loss function for training.
Figure 5: 7-skip frame interpolation visualization results in Adobe240 dataset [56]. Our method (c) more effectively captures the rotating wheels compared to the VideoINR (b), which tends to show noticeable holes and gaps. Green circles highlight obvious holes and gaps.
@@ 4 Experiments
To facilitate a comprehensive comparison between the frame-base [10, 8] and event-guided methods [38, 62, 61, 23], we employed two simulated datasets [56, 40] and two real-world datasets [61, 51] for our experiments (1) Adobe240 dataset [56] includes 133 videos, each with $720\times 1280$ resolution and a $240fps$ frame rate. Following established protocols [10, 8, 70], the dataset is divided into 100 training, 16 validation, and 17 testing subsets. We employed the widely-used event simulation method vid2e [15], which accounts for real noise distribution, enhancing our model’s robustness and generalization capabilities. (2) GoPro dataset [40] featuring the same resolution and frame-rate with Adobe240 dataset, comprises 22 training and 11 test videos. Given its compact size, previous studies [8, 10] have primarily employed the test set for quantitative analysis. We follow this to be consistent with established practice. For datasets (1) and (2) during 7-skip frame interpolation, we input 4 frames where each pair of adjacent frames is separated by 7 frames, allowing a total of 25 frames to serve as GTs. (3) BS-ERGB [61] recorded using a spectroscope, includes real-world paired events and frames. Previous studies, e.g., TimeLens++[61], initially pre-trained on the Vime90k simulation dataset[72] before fine-tuning on BS-ERGB. In contrast, we opted to pre-train our model on the Adobe240 dataset before fine-tuning. Notably, the Vime90k dataset is larger in scale than the Adobe240 dataset. During fine-tuning, we also used perceptual loss [25] with weight $0.1$ to be consistent with previous methods [62, 61] for fair comparison. (4) CED [51] is a real-world dataset in VSR. To fairly compare the previous research [23, 38], only this data set is used for training for VSR comparison.
Implementation Details: Our model is trained using Pytorch [48], employing the Adam optimizer [29]. Referring to the VideoINR [10], our training consists of two stages. (1) Train frame interpolation under a fixed spatial up-sampling (4×4$\times$4 ×), over 70 epochs, starting with a learning rate of $45e-$ 4. (2) Train frame interpolation under random space upsampling rate in $\mathcal{U}(1,8)$, spanning 30 epochs with the learning rate $55e-$ 5. We randomly choose 20 frames from a pool of 25 frames as ground truth. Data augmentation is implemented via Random Crop, extracting $512\times 512$ areas from frames and events, with the input resolution dynamically determined by a randomly chosen upsampling ratio $s$. To optimize memory usage and accelerate speed, we implemented the mixed precision strategy [39, 12]. All experiments are performed on an NVIDIA A800 computing card. Please refer to the Suppl. Mat. for more details.
Evaluation: To ensure the fairness, we adopted PSNR [78], SSIM [67], and LPIPS [78] as quantitative evaluation metrics. Aligning with prior works [10, 8] for consistency, we use only the $Y$-channel for GoPro and Adobe datasets and all three RGB channels for BS-ERGB and CED datasets.
Table 1: Quantitative metrics (PSNR/SSIM) with 7-skip VFI and 4×4$\times$4 × VSR. Center Average remain consistent with previous work [10, 8].
@@
Su-SloMo [22] EDVR [66] 19.8+20.7 28.24/0.8322 26.30/0.7960 27.25/0.7972 25.95/0.7682
BasicVSR [5] 19.8+6.3 28.23/0.8308 26.36/0.7977 27.28/0.7961 25.94/0.7679
@@
DAIN [3] EDVR [66] 24.0+20.7 28.01/0.8239 26.37/0.7964 27.06/0.7895 26.01/0.7703
BasicVSR [5] 24.0+6.3 28.00/0.8227 26.46/0.7966 27.07/0.7890 26.23/0.7725
Zooming Slow Mo [69] 11.10 30.69/0.8847 -/- 30.26/0.8821 -/-
TMNet [70] 12.26 30.14/0.8692 28.83/0.8514 29.41/0.8524 28.30/0.8354
Video INR-fixed [10] 11.31 30.73/0.8850 -/- 30.21/0.8805 -/-
Video INR [10] 11.31 30.26/0.8792 29.41/0.8669 29.92/0.8746 29.27/0.8651
MoTIF [8] 12.55 31.04/0.8877 30.04/0.8773 30.63/0.8839 29.82/0.8750
Table 2: More quantitative comparisons using PSNR/SSIM on the GoPro dataset.
Temporal Scale Spatial Scale Su-SloMo [22] + LIIF [9] DAIN [3] + LIIF [9] TMNet [70] Video INR [10] MoTIF [8] Ours
12$\times$ 4$\times$ 25.07/0.7491 25.14/0.7497 26.38/0.7931 27.32/0.8141 27.77/0.8230 28.87/0.8854
12$\times$ 6$\times$ 22.91/0.6783 22.92/0.6785 - 24.68/0.7358 26.78/0.7908 27.14/0.8173
16$\times$ 4$\times$ 24.42/0.7296 24.20/0.7244 24.72/0.7526 25.81/0.7739 25.98/0.7758 27.29/0.8556
16$\times$ 6$\times$ 23.28/0.6883 $2$ 23.86/0.7123 25.34/0.7527 26.09/0.7954
6$\times$ 1$\times$ - - - 32.34/0.9545 34.77/0.9696 38.53/0.9735

### v2

@@ 3.1 Temporal Pyramid Representation
Firstly, we represent the event stream $\bm{E}$ into a frame-like form that the network can process. To capture holistic motion, we partition all events during $[0,1]$ of the timeline into $M$ equal intervals using a voxel grid (Tulyakov et al, 2021; Gallego et al, 2020), denotes as $\bm{E}_{v}$ with dimensions $M\times H\times W$. In practice, event representation methods like voxel grid (Tulyakov et al, 2021; Gallego et al, 2020) and its extended structure, symmetric cumulative event representation (Sun et al, 2022), achieve time granularity by uniformly dividing time into intervals with resolutions of $1/M$.
@@
Our idea: The core of TPR is constructing a temporal pyramid where each successive layer’s duration is $1/r,(r>1)$ of the preceding one, leading to exponentially finer time granularity with additional layers. For instance, as illustrated in Fig. 2 (a), around any given time $t$, we define a surrounding time window $\Delta t$ and select an attenuation factor $r$. At the pyramid’s $L$-th level, the events are within the time span of $t-\Delta t/r^{L},t+\Delta t/r^{L}]$. Each layer is further segmented into $M_{p}$ intervals, represented using a voxel grid. Accordingly, for an $L$-th layer, each layer contains $M_{p}$ moments within the TPR, and its finest time granularity, denoted by $\delta_{t}$, is as delineated in the Eq. 4. Therefore, for any time $t$, we construct the corresponding TPR $E_{p}$ with shape $L\times M_{p}\times H\times W$. A simplified example is shown in Fig. 3. We record the TPRs at all target timestamp as $\bm{E_{p}}=\{E_{p}\}_{i}^{N_{out}}$.
@@
Discussion: The time granularity $\delta_{t}$ of TPR exponentially improves with the increase in layers - $L$. For an attenuation factor of $r=3$ and a goal to detect motions as brief as $1/1000$ of a second within a $1s$ window ($2\times\Delta t=1$), we require only $7$ TPR layers with each layer divided into $2$ intervals ($M_{p}=2$). Consequently, a TPR with dimensions $7\times 2\times H\times W$ suffices to discern motion down to $1/1000s$. Based on the above representations, we obtained $\bm{E}_{v}$, encapsulating holistic motion, and $\bm{E}_{p}$, which focuses on regional edges and motion.
@@ 3.2 Holistic-Regional Feature Extraction
This module aims to extract features from regional TPRs $\bm{E_{p}}$, and the frame $\bm{I_{in}}$ and the holistic events $\bm{E_{v}}$ for INR-based spatial-temporal decoding. Accordingly, we introduce: (1) the regional event feature extractor $f_{re}$, Eq. LABEL:eq:feature_extraction (a), for dynamic motion and edges detail capture. (2) the holistic event-frame feature extractor $f_{he}$, Eq. LABEL:eq:feature_extraction (b), for long-term motion dependencies modeling across time and space.
@@
Figure 4: Holistic event-frame feature extractor. The down-sample module will halve the resolution. The up-sample module will double the resolution. The encoder and decoder have the same structure as Swin-Transformer (Liu et al, 2022b, 2021; Geng et al, 2022).
[…]
@@
Figure 5: Temporal embedding. Given the input time $t\in[0,1]$, the output is the temporal attention $E_{t}$ derived from a two-layer MLP. (b) presents a visualization of the trained $E_{t}$ during $[0,1]$ on real-world dataset (Tulyakov et al, 2022).
(1) Regional Event Feature Extraction: The regional event feature extractor $f_{re}$ takes as input the Temporal Pyramid Representation (TPR) $\bm{E_{p}}\in\mathbb{R}^{L\times M_{p}\times H\times W}$ at each timestamp $t$, where $L$ denotes the number of TPR layers and $M_{p}$ represents the number of moments within each layer. Since $f_{re}$ is invoked $N_{out}$ times, its design must balance computational efficiency with the capacity to model inter-level relationships effectively, ensuring the accurate capture of regional motion and edge details. The feature extraction process is divided into two stages: preprocessing and the main feature extraction via Swin Transformer Encoder Blocks (STEB) (Liu et al, 2022b, 2021; Geng et al, 2022). In the preprocessing stage, a $1\times 1$ convolutional layer is applied to increase the feature dimensions to $C_{r}$, resulting in a shape of $L\times C_{r}\times H\times W$. This step enhances the feature’s capacity to capture intricate details from the event stream. In the subsequent feature extraction stage, four STEBs are employed to model the relationships between different pyramid levels. STEB utilizes a multi-head self-attention mechanism, which enables the model to effectively capture long-range dependencies and convey edge information across different levels of the pyramid. This mechanism improves the model’s ability to capture short-term motion while maintaining computational efficiency through parameter optimization. The input and output dimensions of STEB remain consistent throughout the process, preserving the feature shape $L\times C_{r}\times H\times W$. Thus, the $f_{re}$ module incorporates both preprocessing and feature extraction steps to ensure efficient extraction of regional motion features while maintaining high computational efficiency.
Swin Transformer Encoder Blocks (STEB): The input and output shapes of STEB are represented as $L\times C\times H\times W$ for clarity. Initially, for a window size of $M\times M$, specifically $4\times 4$ in our implementation, the input is partitioned into disjoint windows of $(M\times M)\times L\times(H/M)\times(W/M)\times C$ dimensions. Then, each window is compressed to form a feature map of shape $(M\times M)\times(L\times H\times W/M^{2})\times C$. Following this, Layer Normalization (Ba et al, 2016) and window-based multi-head self-attention (Liu et al, 2021; Geng et al, 2022) are computed for each window, succeeded by further transformation via another Layer Normalization and a Multi-Layer Perception. Shifted window-based multi-head self-attention (Liu et al, 2021; Geng et al, 2022) is then employed to establish cross-window connections. After applying one STEB structure, a second STEB is introduced with an identical configuration, except the input feature window is offset by $(M/2)\times(M/2)$. In total, four STEBs are utilized for comprehensive feature extraction in the Regional Event Feature Extraction.
(2) Holistic Event-Frame Feature Extraction: The inputs of $f_{he}$ is $\bm{I_{in}}$ and $\bm{E_{v}}$, as shown in Fig. 4. First, a convolutional layer processes both frames and events to increase dimensions. Motivated by the events feature manifest between successive frames to provide inter-frame motion information. We construct $N_{in}-1$ event segments so that each segment complements the inter-frame motion between two neighboring RGB frames. We then incorporate the STEB to facilitate interactions across varied levels and spatial domains. To minimize computational overhead and expand the receptive field, we integrated a down-sampling module between STEBs, forming a multi-scale encoder. Each down-sampling iteration halves the resolution while maintaining the channel dimensions. After three iterations, the feature resolution reduces to $1/8$ of its initial size, enlarging the receptive field. We then employ a Swin Transformer Decoder Block (Liu et al, 2022b, 2021; Geng et al, 2022) to fuse features at matching resolutions. Each of the first three STEBs is followed by an upsampling process, which doubles the resolution while maintaining the channel count. Ultimately, this process outputs the feature $F_{g}$.
[…]
Figure 6: 7-skip frame interpolation visualization results in Adobe240 dataset (Su et al, 2017). Our method (c) more effectively captures the rotating wheels compared to the VideoINR (b), which tends to show noticeable holes and gaps. Green circles highlight obvious holes and gaps.
@@ 3.3 INR-based Spatial-Temporal Decoding
Spatial Embedding and Decoding: To upscale the temporal embedded features to any desired spatial scale, we utilize a similar approach to previous works (Chen et al, 2022, 2023b, 2021). We query the four nearest neighbors in the temporal embedded feature for each spatial coordinate and concatenate these with their distances for spatial embedding. A four-layer MLP decoder computes the RGB values, which are then aggregated through area-weighted interpolation for arbitrary-scale super-resolution. Similar to (Lu et al, 2023), we use the Charbonnier loss (Lai et al, 2018) as the fundamental loss function for training.
@@ 4 Experiments
In this section, we first introduce the datasets, network training and testing details, as well as the evaluation methods in Sec. 4.1. Then, we conduct comparative experiments, which include both joint C-STVSR comparisons and individual comparisons for VFI and VSR tasks in Sec. 4.2. Furthermore, in the following analysis and ablation studies, we explore the performance and characteristics of various components within our framework, as discussed in Sec. 4.3.
4.1 Datasets and Implementation Setting
Table 1: Quantitative metrics (PSNR/SSIM) with 7-skip VFI and $4\times$ VSR. Center Average remain consistent with previous work (Chen et al, 2022, 2023b).
@@ 4.1 Datasets and Implementation Setting
Su-SloMo
EDVR 19.8+20.7 28.24/0.8322 26.30/0.7960 27.25/0.7972 25.95/0.7682
BasicVSR 19.8+6.3 28.23/0.8308 26.36/0.7977 27.28/0.7961 25.94/0.7679
Bicubic 29.2 26.50/0.7791 25.41/0.7554 25.57/0.7324 24.72/0.7114
QVI
EDVR 29.2+20.7 27.43/0.8081 25.55/0.7739 26.40/0.7692 25.09/0.7406
BasicVSR 29.2+6.3 27.44/0.8070 26.27/0.7955 26.43/0.7682 25.20/0.7421
Bicubic 24.0 26.92/0.7911 26.11/0.7740 26.01/0.7461 25.40/0.7321
DAIN
EDVR 24.0+20.7 28.01/0.8239 26.37/0.7964 27.06/0.7895 26.01/0.7703
BasicVSR 24.0+6.3 28.00/0.8227 26.46/0.7966 27.07/0.7890 26.23/0.7725
TimeLens EG-VSR 72.2+2.45 28.85/0.8678 27.54/0.8293 28.11/0.8441 27.42/0.8269
CBMNet EG-VSR 22.2+2.45 29.22/0.8686 28.51/0.8493 28.28/0.8553 27.89/0.8334
Zooming Slow Mo 11.10 30.69/0.8847 -/- 30.26/0.8821 -/-
TMNet 12.26 30.14/0.8692 28.83/0.8514 29.41/0.8524 28.30/0.8354
Video INR-fixed 11.31 30.73/0.8850 -/- 30.21/0.8805 -/-
Video INR 11.31 30.26/0.8792 29.41/0.8669 29.92/0.8746 29.27/0.8651
MoTIF 12.55 31.04/0.8877 30.04/0.8773 30.63/0.8839 29.82/0.8750
Table 2: More quantitative comparisons using PSNR/SSIM on the GoPro dataset. Bold indicates the best performance.
Temporal
Scale
Spatial
Scale
Su-SloMo
+ LIIF
DAIN
+ LIIF
TMNet Video INR MoTIF Ours
@@
[…]
Datasets: To facilitate a comprehensive comparison between the frame-base (Chen et al, 2022, 2023b) and event-guided methods (Lu et al, 2023; Tulyakov et al, 2021, 2022; Jing et al, 2021), we employed two simulated datasets (Su et al, 2017; Nah et al, 2017) and two real-world datasets (Tulyakov et al, 2022; Scheerlinck et al, 2019a) for our experiments. (1) Adobe240 Dataset (Su et al, 2017): This dataset comprises 133 videos, each with a resolution of $720\times 1280$ and a frame rate of 240. We follow (Chen et al, 2022, 2023b; Xu et al, 2021) to split this dataset into 100 training, 16 validation, and 17 test sets. We employed the widely-used event simulation method vid2e (Gehrig et al, 2020), which accounts for real noise distribution, ensuring robust neural network training with enhanced generalization. In generating the inputs and ground truth of the network, we adopted and extended the previous works (Chen et al, 2023b, 2022) to accommodate multi-frame input. Specifically, input and ground truth (GT) frames are selected via sliding windows. We define the window size as $W$, the number of input frames as $N_{in}$, and the interval as $S$. They interrelate as: $W=(N_{in}-1)*(S+1)+1$ For instance, with 4 frames input at 7-frame intervals, 25 frames are chosen. The 1st, 9th, 17th, and 25th frames become the input after down-sampling, while frames 1-25 serve as the GT. We adopted two strategies in line with prior works (Chen et al, 2022, 2023b): I) A fixed magnification set at $4\times$ the input resolution, and II) A variable enlargement strategy, wherein the scaling factor is governed by a $\mathcal{U}(1,8)$ distribution. (2) GoPro dataset (Nah et al, 2017) featuring the same resolution and frame-rate with  Adobe240 dataset, comprises 22 training and 11 test videos. Given its compact size, previous studies (Chen et al, 2023b, 2022) have primarily employed the test set for quantitative analysis. We follow this to be consistent with established practice. (3) BS-ERGB (Tulyakov et al, 2022) recorded using a spectroscope, includes real-world paired events and frames with a $970\times 625$ resolution at 28 fps. Of the 123 videos in the dataset, 47 are allocated for training, 19 for validation, and 26 for testing. Each video contains between 100 to 600 frames. We opted to pre-train our model on the Adobe240 dataset before fine-tuning. During fine-tuning, we also used perceptual loss (Johnson et al, 2016) with weight $0.1$ to be consistent with previous methods (Tulyakov et al, 2021, 2022) for fair comparison. (4) CED (Scheerlinck et al, 2019a) is a real-world dataset in VSR. To fairly compare the previous research (Jing et al, 2021; Lu et al, 2023), only this data set is used for training for VSR comparison. It is worth noting that the dataset ALPIX-VSR (Lu et al, 2023) can also be used for event-guided VSR; however, its events are not stream-based  (Alpsentek, 2024), making it unsuitable for events TPR. Therefore, we use CED as the primary dataset for comparison.
Figure 7: Timestamps and corresponding PSNR for each frame during $4\times$ VSR and 7-skip VFI on the Adobe240 dataset (Su et al, 2017).
Implementation Details: Our model is trained using Pytorch (Paszke et al, 2019), employing the Adam optimizer (Kingma and Ba, 2014). Referring to the VideoINR (Chen et al, 2022), our training consists of two stages. (1) Train frame interpolation under a fixed spatial up-sampling ($4\times$), over 70 epochs, starting with a learning rate of 5e45e-4. (2) Train frame interpolation under random space upsampling rate in $\mathcal{U}(1,8)$, spanning 30 epochs with the learning rate 5e55e-5. We randomly choose 20 frames from a pool of 25 frames as ground truth. Data augmentation is implemented via Random Crop, extracting $512\times 512$ areas from frames and events, with the input resolution dynamically determined by a randomly chosen upsampling ratio $s$. To optimize memory usage and accelerate speed, we implemented the mixed precision strategy (Micikevicius et al, 2017; Das et al, 2018). All experiments are performed on an NVIDIA A800.
Figure 8: 3-skip frame interpolation visualization results in BS-ERGB dataset (Tulyakov et al, 2022). Our method accurately captures local non-linear motion (e.g., balls in (1.d) and (2.d)), surpassing TimeLens (Tulyakov et al, 2022), which exhibits ghosting and holes (green circles). A yellow arrow shows TimeLens’s inaccurate ball positioning.
Evaluation: To ensure the fairness, we adopted PSNR (Zhang et al, 2018), SSIM (Wang et al, 2004), and LPIPS (Zhang et al, 2018) as quantitative evaluation metrics. For a fair comparison on GoPro and Adobe datasets, we follow previous work (Chen et al, 2022, 2023b) by using only the $Y$-channel, while for the BS-ERGB and CED datasets, we employ all three RGB channels.